A property has a market value of $500,000. The taxing authority taxes 70% of the value and the property has a homestead exemption of $50,000. The county, town and board of education have tax rates of 5.8, 6.9 and 2.3, respectively. What are the taxes for the property?
Answer #1

Market Value of the property = $500,000

Assessment ratio = 70%

Assessed value is $500,000*70% = $350,000

Exemption available $50,000  

Taxable value = 350,000-50,000 = $300,000.

Taxes for the property 300,000 * 5.8% = $17400 Country

300,000 * 6.9% =$ 20,700 town

300,000 * 2.3% = $6900 School

Hope that helped.Do share your feedback Thank you